Please help me solve this word problem:
The numbers A and B each have 3 digits. Zoe was asked to calculate AxB. Instead she put A on the left of B to form 6 digit numbers, D. Her answer was 3 times the correct answer. What is A and B?
_______
I have thought of reversing the orders, trying to find D and then i would figure out how to find out what A and B is...
Please help me, and your help will be very much appreciated!

Answer by ankor@dixie-net.com(22740)

The numbers A and B each have 3 digits. Zoe was asked to calculate AxB. Instead she put A on the left of B to form 6 digit numbers, D. Her answer was 3 times the correct answer. What is A and B?
:
For what it's worth, (not much you will probably say)
I can show you how to solve it using a graphing calc like a Ti83
:
The problems states:
1000A + B = 3(AB)
:
B - 3AB = - 1000A
B(1-3A) = - 1000A
B = -1000A/(1-3A)
:
enter into calc: y = -1000x/(1-3x): use the x,y table
The first integers that satisfy the requirements are:
A = 167, B= 334
3(167*334) = 167334